# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

### v9

@@ 4 NDGM and the LASSO penalty
Our experiments yielded two key observations (Figure 3): 1) The $L_{1}$ norms at convergence differ substantially across the three optimizer variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence. These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may be more complex than previously understood. Indeed, our results indicate that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 2): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, more than 90% do. Furthermore, these numbers are sensitive to the weight initialization bringing into question the notion of “model interpretability" in the presence of non-differentiability. These observations raise important questions about the efficacy and generalizability of $L_{1}$ regularization for network pruning across different optimization algorithms and regularization strengths.
@@ A.2 Code for activation function comparison experiments
A.3 A Geometric Characterization of Strong Duality
@@ A.3.2 Geometric Properties of Global Minima
In our specific case, since $f$ has a global minimum at 0, $w^{*}$ represents the global lowest point on the epigraph an represents the lowest point of the epigraph along the $(n+1)$st axis. Our proofs utilize the following result:
@@
We first show that $(0,\ldots,0,w^{*})$ is not an interior point of $M$. To see why, suppose by contradiction that it is. Then there exists $\epsilon>0$ such that $B_{\epsilon}((0,\ldots,0,w^{*}))\subset M$. This implies $(0,\ldots,0,w^{*}-\epsilon/2)\in M$, contradicting that $w^{*}$ is the value at the global minimum.
@@
To prove non-verticality, suppose by contradiction that some $h\in\mathcal{H}$ is vertical. Since $h$ passes through $(0,\ldots,0,w^{*})$, it must have the form $q^{T}\beta=0$ for some $q\in\mathbb{R}^{n}$, where $q\neq 0$. Construct $u\in\mathbb{R}^{n}$ by setting $u[i]=\text{sign}(q[i])$ for each component. Then $q^{T}u>0$ by construction. Note that points $(u,f(u))$ and $(-u,f(-u))$ must belong to $M$, where $f(u)$ and $f(-u)$ represent the height of the epigraph at these points. The heights at these points can be $\infty$ but not $\infty$- ∞ as this would violate the existence of a finite global minimum. Importantly, the actual values of these heights do not affect our argument. Since $M$ must lie in the upper half-space of $h$, we require $q^{T}(-u)>0$. However, $q^{T}(-u)=-(q^{T}u)<0$, a contradiction. ∎
@@
A standard result in convex analysis is that all closed proper convex functions show strong duality (see Section 4.2.1 in Bertsekas (2009) for a demonstration of this fact using the MC/MC framework). Our conditions are at least as strong as this condition as shown in the following proposition:
@@
Any closed proper convex function satisfies the conditions shown above [^10] It is possible that the stronger ”if and only if” condition holds, but we are unable to show the reverse direction, and leave it for future work.
@@
The above relationship connects naturally to conjugate function theory. As shown in Section 4.2.1 of Bertsekas (2009), for the MC/MC framework we have $w^{*}=f(0)$ and $q^{*}=f^{**}(0)$, providing another perspective on our result: for closed proper convex functions, the equality $q^{*}=w^{*}$ aligns with the fundamental property that $f(x)=f^{**}(x)$ for all $x\in\mathbb{R}^{n}$. By enforcing some mild constraints on the nature of $f$, we can provide an even stronger qualification of the dual optimal solution as is shown by the proposition below
@@
Now, for the reverse direction, suppose that the set of solutions for the max crossing problem are non-empty and compact. Then, from Proposition 4.4.2 of Bertsekas (2009), the origin has to be an interior point of $D$ described in point number 3 above. This would require 0 to be an interior point of $dom(f)=\left\{x\in X\text{ | }f(x)<\infty\right\}$. Since $f$ has a finite global minimum at 0, we know that 1) $f$ is proper, and 2) $0\in dom(f)$. To establish a contradiction, suppose $f$ is discontinuous at 0. Then, 0 will not be in the relative interior of $dom(f)$ (if $f$ is a proper convex function on (,]($\infty$,$\infty$]( - ∞ , ∞ ], then $f$ restricted to $dom(f)$ is continuous over the relative interior of $dom(f)$ (Proposition 1.3.11 in Bertsekas (2009))), which would imply that 0 is not in the interior of $dom(f)$, which would imply that the solutions to the max crossing problem are not compact, thus completing the proof.
@@
Special case of the previous proposition ∎
@@
This result is remarkably general - it requires nothing about the function beyond the existence of a finite value at the minimum. The geometry alone precludes vertical supporting hyperplanes.
@@ A.3.4 A plausible explanation for the non-convergence of the ReLU networks with the Capture Theorem
$=\arg\min_{\beta}\left\|\mathbf{y}_{1}-\max(0,\mathbf{Z}_{1}\beta)\right\|^{2}$
@@
In the extended value formulation, $Q(\beta)$ takes a finite value only at $\beta=0$, and is infinite elsewhere, demonstrating the equivalence to 11. The Epigraph of $Q(\beta)$ is given by
@@
Appendix B Appendices for section 4
@@ B.3.4 Adam and the LASSO
From 29, we observe that $m_{t+1}[k]$ is a convex combination of $m_{t}[k]$ and either $-\lambda_{1}$ or $\lambda_{1}$. By induction, if $|m_{t}[k]|<\lambda_{1}$, then $|m_{t+1}[k]|<\lambda_{1}$. Since $m_{0}[k]=0$, the result follows. ∎
@@
Suppose $m_{T}[k]=$. Using $m_{0}[k]=0$, we can write 29 as
@@
For our proof we use $\alpha_{t}=\frac{\alpha}{\sqrt{v_{t}[k]+\bar{\epsilon}}}$, noting that $\alpha_{t}>0$ for all $t>0$, and is independent of $\beta_{0}$ and $\delta$. Suppose the $k^{th}$ entry in the vector becomes 0 after $N+1$ steps. Then, completing the recursion in 30 using (31) with $m_{0}[k]=0$ gives
@@
Since $\beta_{N}[k]=0$, we have $\nabla L_{3}(\beta_{N}[k])=0$, and therefore, from 29, $m_{N+1}[k]=\delta m_{N}[k]$. The result follows by substituting $m_{N+1}[k]$ into 30. ∎
@@
Using 29 and the definition of a run, we have

### v11

@@ 4 NDGM and the LASSO penalty
Our experiments yielded two key observations (Figure 3): 1) The $L_{1}$ norms at convergence differ substantially across the three optimizer variants, and 2) for SGD and Adam, larger values of $\lambda_{1}$ result in a larger $L_{1}$ norm at convergence. These findings suggest that the relationship between $L_{1}$ regularization and prunable parameters may be more complex than previously understood. Indeed, our results indicate that a larger $\lambda_{1}$ does not necessarily correspond to more prunable parameters (Table 2): with $\lambda_{1}=2$, less than 0.2% of the parameters at convergence with vanilla NDGM have absolute value less than $10^{-5}$, but with $\lambda_{1}=0.002$, more than 90% do. Furthermore, these numbers are sensitive to the weight initialization and batch order during optimization bringing into question the notion of “model interpretability" in the presence of stochasticity and non-differentiability. These observations raise important questions about the efficacy and generalizability of $L_{1}$ regularization for network pruning across different optimization algorithms and regularization strengths.
@@ 5 NDGM and the Edge of Stability
6 Painful Pathologies of Poorly Programmed Perturbations
 [^9] The authors had initially named this section ”LLMs and Training”, but deemed it to be way too boring. Woot! Got an alliteration for yo ass!
Training of modern LLMs often follows schedules derived from empirical scaling laws (Kaplan et al., 2020) and compute-optimal insights such as those in Hoffmann et al. (2022). These protocols typically advocate stopping well before full convergence—arguing that beyond a certain point, the returns diminish while compute costs dominate. The prevailing view is that compute, not convergence, is the primary bottleneck, and that training all the way to the global minimum is neither feasible nor necessary.
But even if the minimum remains out of reach, non-differentiability cannot be ignored. The breakdown of differentiability—and with it, the erosion of duality and geometric support—is not an asymptotic fluke. It is a structural inevitability in wide, deep, non-smooth architectures. As training descends deeper into the loss landscape, the geometry turns hostile: support vanishes, non-verticality collapses, and oscillations emerge—not due to bad hyperparameters, but because the landscape itself no longer admits reliable directional information.
This makes it imperative to analyze the dynamics of modern training under non-differentiability and finite step sizes—not as pathological edge cases, but as central features of the large-scale regime.
@@ A.2 Code for activation function comparison experiments
A.3 Dense Digressions into Devious Duality
@@ A.3.2 Geometric Properties of Global Minima
In our specific case, since $f$ has a global minimum at 0, $w^{*}$ represents the global lowest point on the epigraph and represents the lowest point of the epigraph along the $(n+1)$st axis. Our proofs utilize the following result:
@@
We are aware of three ways of establishing this result: one geometrical, and two analytical. The first two approaches rely on the fact that $(0,\ldots,0,w^{*})$ is not an interior point of $M$, so we begin by establishing this claim before presenting the main argument from all three perspectives.
To show that $(0,\ldots,0,w^{*})$ is not an interior point of $M$, suppose by contradiction that it is. Then there exists $\epsilon>0$ such that $B_{\epsilon}((0,\ldots,0,w^{*}))\subset M$. This implies $(0,\ldots,0,w^{*}-\epsilon/2)\in M$, contradicting that $w^{*}$ is the value at the global minimum.
The geometric approach:
@@
To prove non-verticality, suppose by contradiction that some $h\in\mathcal{H}$ is vertical. Since $h$ passes through $(0,\ldots,0,w^{*})$, it must have the form $q^{T}\beta=0$ for some $q\in\mathbb{R}^{n}$, where $q\neq 0$. Construct $u\in\mathbb{R}^{n}$ by setting $u[i]=\text{sign}(q[i])$ for each component. Then $q^{T}u>0$ by construction. Note that points $(u,f(u))$ and $(-u,f(-u))$ must belong to $M$, where $f(u)$ and $f(-u)$ represent the height of the epigraph at these points. The heights at these points can be $\infty$ but not $\infty$- ∞ as this would violate the existence of a finite global minimum. Importantly, the actual values of these heights do not affect our argument. Since $M$ must lie in the upper half-space of $h$, we require $q^{T}(-u)>0$. However, $q^{T}(-u)=-(q^{T}u)<0$, a contradiction.
Analytical approach 1 - Using the non-vertical hyperplane theorem:
Since 1) $(0,\ldots,0,w^{*})$ is not in the interior of $M$, and 2) $M$ is a convex set with no vertical lines, the non-vertical hyperplane theorem (Proposition 1.5.8 in Bertsekas (2009)) states that there must exist a non-vertical hyperplane passing through $(0,\ldots,0,w^{*})$ that contains $M$ in its positive halfspace. Furthermore, by the definition of max-crossing, the highest intercept on the $(n+1)st$ axis of any hyperplane containing $M$ in its positive halfspace is $q^{*}$. Therefore, for any $(u,w)\in M$, there exists some $\mu\in R^{N}$, such that
$\mu^{T}u+w\geq q^{*}\geq w^{*}.$ (11)
However, by weak duality (Proposition 4.1.2 in Bertsekas (2009)) we have $w^{*}\geq q^{*}$, so combining the two inequalities gives $q^{*}=w^{*}.$
Analytical approach 2 - Using the non-vertical hyperplane theorem and the sandwich theorem:
To derive the result a third way, we first show that $(0,w^{*}-\epsilon)\notin cl(M)$ for any $\epsilon>0$. To see why, suppose by contradiction that it is. Then, there would exist a sequence $(u_{k},w_{k})\in M$ such that $u_{k}\rightarrow 0$ and $w_{k}\rightarrow w^{*}-\epsilon$, which would imply that there exists an $N_{0}$ such that $w_{N}<w^{*}$ $\forall N\geq N_{0}$. This is not possible since $w^{*}$ is the global minimum, thus establishing the contradiction.
Accordingly, by the non-vertical hyperplane theorem (Proposition 1.5.8b in Bertsekas (2009)), there exists a non vertical hyperplane that strictly separates $(0,w^{*}-\epsilon)$ and $M$. Consider such a hyperplane which has an intercept of $\xi$ on the $(n+1)^{st}$ axis, and contains $(0,w^{*}-\epsilon)$ in the negative halfspace, and $M$ in the positive halfspace. Since $q^{*}$ is the highest intercept of a non-vertical hyperplane having $M$ in its positive halfspace, we have $\xi\leq q^{*}$, and by weak duality we have $q^{*}\leq w^{*}$, so putting it altogether gives
$w^{*}-\epsilon<\xi\leq q^{*}\leq w^{*}.$ (12)
The result follows as $\epsilon\rightarrow 0$. ∎
What a difference finiteness makes!:
The result above has strong connections with the non-vertical hyperplane theorem stated below for completeness
Proposition A.5.
Let $C$ be a non empty convex subset of $R^{N+1}$ that contains no vertical lines. Let the vector in $R^{N+1}$ be denoted by $(u,w)$, where $u\in R^{N}$ and $w\in R$. Then $C$ is contained in a closed halfspace corresponding to a non-vertical hyperplane.
Proof.
Simplified version of Proposition 1.5.8 in Bertsekas (2009). ∎
Note that the above proposition is non-comittal on whether the non-vertical hyperplane “touches" $C$ or not. Our proposition shows that “touching" is guaranteed if the global minimum is finite, thus providing the bridge between weak and strong duality. This point is alluded to indirectly via the max-crossing problem, but is not explicitly stated to the best of our knowledge.
Geometric intuition:
@@
A standard result in convex analysis is that all closed proper convex functions show strong duality (see Section 4.2.1 in Bertsekas (2009) for a demonstration of this fact using the MC/MC framework). Our conditions are stronger than this condition as shown in the following proposition:
@@
Any closed proper convex function satisfies the conditions shown above.
@@
To demonstrate that our result is indeed stronger, consider the function
$\widehat{f(x)}=\begin{cases}0&\text{ if $x\in(-1,1)$}\\ \infty&\text{otherwise},\end{cases}.$ (13)
This function above satisfies the conditions above, but is not closed. The reason our result is stronger is because lower semi-continuity is a local property of the function, whereas closedness is a global property of the function. More precisely, closedness implies lower semi-continuity at every point in the domain (Proposition 1.1.2 in Bertsekas (2009)). However as we show in Proposition A.3, our results require lower semi-continuity at only one point – 0.
Connection to conjugate function theory:
The above proposition connects naturally to conjugate function theory. As shown in Section 4.2.1 of Bertsekas (2009), for the MC/MC framework we have $w^{*}=f(0)$ and $q^{*}=f^{**}(0)$, providing another perspective on our result: for closed proper convex functions, the equality $q^{*}=w^{*}$ aligns with the fundamental property that $f(x)=f^{**}(x)$ for all $x\in\mathbb{R}^{n}$. By enforcing some mild constraints on the nature of $f$, we can provide an even stronger qualification of the dual optimal solution as is shown by the proposition below
@@
Now, for the reverse direction, suppose that the set of solutions for the max crossing problem are non-empty and compact. Then, from Proposition 4.4.2 of Bertsekas (2009), the origin has to be an interior point of $D$ described in point number 3 above. This would require 0 to be an interior point of $dom(f)=\left\{x\in X\text{ | }f(x)<\infty\right\}$. Since $f$ has a finite global minimum at 0, we know that 1) $f$ is proper, and 2) $0\in dom(f)$. To establish a contradiction, suppose $f$ is discontinuous at 0. The discontinuity implies that the function has to “jump" at 0 to a value different than $f(0)$. There are only two possibilities:
  • The function jumps to an infinite value at 0 - This is not possible because 0 would not be an interior point of $dom(f)$, which is required for compactness.
  • The function jumps to a finite value different than $f(0)$ This is not possible because in this case, 0 will not be in the relative interior of $dom(f)$ (if $f$ is a proper convex function on (,]($\infty$,$\infty$]( - ∞ , ∞ ], then $f$ restricted to $dom(f)$ is continuous over the relative interior of $dom(f)$ (Proposition 1.3.11 in Bertsekas (2009))), which would imply that 0 is not in the interior of $dom(f)$, which would imply that the solutions to the max crossing problem are not compact
The result follows since neither possibility is feasible.
@@
Special case of the previous proposition. ∎
@@
This result requires nothing about the function beyond the existence of a finite value at the minimum. The geometry alone precludes vertical supporting hyperplanes. We can generalize the logic even further to incorporate any function whose domain is $R^{N}$ as seen in the proposition below.
Proposition A.12.
[…]
Proof.
As we discussed in the geomtric proof of Proposition A.4, the non-existence of non-vertical supporting hyperplane at a point is only predicated on the existence of symmetric points on either side of that point in the domain. The result follows since the domain of $f$ is $R^{N}$, and $u$ is finite. ∎
Geometry in action:
To exemplify the generality of our geometric approach, we “work through an example" using a variant of the Dirichlet function which is of historical importance. The function is defined as
$f_{d}(x)=\begin{cases}0&\text{ if x is rational}\\ 1&\text{otherwise},\end{cases}.$ (14)
This is the “worst case scenario" we could conceive of for Proposition A.11 – a function with a global minimum at 0 with no continuity, differentiability or convexity to hang its coat on. The proposition claims that a non vertical supporting hyperplane supporting $Epi(f_{d})$ at 0 will not exist if there are points on either side of 0 belonging to the domain, a condition which is clearly being satisfied – for any finite $u$, $f_{d}(u)$ and $f_{d}(-u)$ exist. This claim is easily verified for this 1-D case – a vertical hyperplane on $Epi(f_{d})$ at 0 would split the epigraph into two symmetric halves, with the positive and negative values of $x$ lying on different sides of the hyperplane. Similarly, Proposition A.12 guarantees that the “vanilla" Dirichlet function
$f^{\prime}_{d}(x)=\begin{cases}1&\text{ if x is rational}\\ 0&\text{otherwise},\end{cases}.$ (15)
cannot have a non-vertical hyperplane at 0 containing $Epi(f^{\prime}_{d})$ in its positive halfspace.
@@ A.3.4 A plausible explanation for the non-convergence of the ReLU networks with the Capture Theorem
$=\arg\min_{\beta}\left\|\mathbf{y}_{1}-\max(0,\mathbf{Z}_{1}\beta)\right\|^{2}+a_{1}\|\beta\|_{1}+a_{2}\|\beta\|_{2}^{2},$
@@
In the extended value formulation, $Q(\beta)$ takes a finite value only at $\beta=0$, and is infinite elsewhere, demonstrating the equivalence to 16. The Epigraph of $Q(\beta)$ is given by
@@
A moment of reflection:
While it is plausible that some of the mathematics in this appendix is useful, it contains a hidden gem of undeniable value. The appendix title includes the second alliteration! We are now on rarefied ground.
Appendix B An Axiomatic Anthology of Numerical Alliterations
@@ B.3.4 Adam and the LASSO
From 34, we observe that $m_{t+1}[k]$ is a convex combination of $m_{t}[k]$ and either $-\lambda_{1}$ or $\lambda_{1}$. By induction, if $|m_{t}[k]|<\lambda_{1}$, then $|m_{t+1}[k]|<\lambda_{1}$. Since $m_{0}[k]=0$, the result follows. ∎
@@
Suppose $m_{T}[k]=$. Using $m_{0}[k]=0$, we can write 34 as
@@
For our proof we use $\alpha_{t}=\frac{\alpha}{\sqrt{v_{t}[k]+\bar{\epsilon}}}$, noting that $\alpha_{t}>0$ for all $t>0$, and is independent of $\beta_{0}$ and $\delta$. Suppose the $k^{th}$ entry in the vector becomes 0 after $N+1$ steps. Then, completing the recursion in 35 using (36) with $m_{0}[k]=0$ gives
@@
Since $\beta_{N}[k]=0$, we have $\nabla L_{3}(\beta_{N}[k])=0$, and therefore, from 34, $m_{N+1}[k]=\delta m_{N}[k]$. The result follows by substituting $m_{N+1}[k]$ into 35. ∎
@@
Using 34 and the definition of a run, we have
@@ B.4 Connections to deep learning training theory:
A moment of reflection:
It goes without saying that the title of this appendix contains the third alliteration. Three strikes and you’re out! Wait, what? You didn’t even notice? WTF bro? Are you even paying attention? Haven’t you read the gospel of Vaswani? – Attention is all you need! You should totally read the paper one more time.  [^11] I know, the “Numerical” kinda ruins the vibe a bit, but didn’t want to cheat and use GPT for this shit. Sigh! If I had only listened to Tulja and studied Strunk and White more closely :(. But, in my defense – “omit needless words”? Come on now! That’s not sage advice to any a stoner. Gotta have your own swag baby! To quote the great Calvin Harris - “My way! Oh way, oh way, oh way”.